# State-space kinetic Ising model reveals task-dependent entropy flow in sparsely active nonequilibrium neuronal dynamics

### v2

@@ Results
The term $r\left(g_{i}+z\sqrt{\Delta_{i}}\right)-m_{i}$ quantifies how the neuron’s activity rate deviates from its long-term average, while $g_{i}+z\sqrt{\Delta_{i}}$ represents the fluctuating input it receives. The steady-state mean-field solution thus provides an intuitive view of entropy flow as a measure of a neuron’s causal responsiveness to input fluctuations – a quantity that captures the correlation underlying Hebbian plasticity in neural systems. However, this equation also clarifies that the approximation depends mainly on the magnitudes of the field and coupling parameters and remains insensitive to the detailed coupling structure. It should therefore be applied with caution when asymmetry is the primary source of entropy flow.
Application to simulated data
Estimating the model parameters from simulated data
[…]
@@
We start testing the proposed method by estimating the time-dependent parameters of a kinetic Ising model consisting of two simulated neurons (Fig. 1A). Figure 1B shows the spike data generated using Eq. (1) with the number of bins, $T=400$, and the number of trials, $R=200$. The time-dependent parameter ${\boldsymbol{\theta}}_{t}$ used in the data generation process was sampled from Gaussian processes (See Methods).
@@
The EM algorithm, as described in Methods, was applied to this spike data until the log marginal likelihood converges (Fig. 1C). Figure 1D shows the components of the optimized hyperparameter matrices, $\mathbf{Q}^{i}(i=1,2)$. Figures 1E show the MAP estimates of the time-dependent fields $\theta_{i,t}$ and interactions $\theta_{ij,t}$ under the optimized hyperparameters (solid lines) with 95% credible intervals (shaded areas). The results confirm that the method uncovers the underlying time-dependent parameters (black dashed lines) used to generate the data.
@@
Fig 2: The application of the state-space kinetic Ising model to 12 simulated neurons. A Simulated spike data for the first, 100th, and last trial out of 200 trials. The vertical axis shows the number of neurons, and the horizontal axis represents the number of bins. B Estimated coupling parameters $\boldsymbol{\theta}_{t|T}$, for all neurons and time bins ($i=1,2,\ldots,12$, $t=1,\ldots,T$). C Scatter plots comparing the true coupling parameters ${\boldsymbol{\theta}}_{t}$ with the estimated values $\boldsymbol{\theta}_{t|T}$ at time $t=10,20,\ldots,60$. The black line is a diagonal line.
@@
Next, we applied the state-space kinetic Ising model to a network of 12 simulated neurons to estimate the time-varying field and coupling parameters between neurons. Figure 2A presents the spike data generated using the observation model with the number of bins set to $T=75$ and the number of trials $R=200$. Data simulation and estimation procedures follow the two-neuron case above. Figure 2B shows the estimated time-varying coupling parameters $\theta_{ij,t}$ for each neuron. These plots show only the couplings that are significantly deviated from zero: The couplings whose 95% credible interval contain 0 in all bins were excluded. In Fig. 2C, we compare the estimated coupling parameters $\boldsymbol{\theta}_{t|T}$ with the true values $\boldsymbol{\theta}_{t}$ at representative time points ($t=10,20,\ldots,60$). The scatter plot shows agreement between the true and estimated values, with most points aligning closely along the diagonal line, indicating that the model captured the underlying dynamics of the coupling parameters. These results confirm that the proposed state-space kinetic Ising model can reliably estimate time-varying coupling parameters in a network of simulated neurons.
Estimation error and computational time
@@
Fig 3: Estimation error and computational time. A Root mean squared error (RMSE) for field and coupling parameters as a function of trials $R$, with the number of neurons fixed at $N=80$. The results are averaged over 10 independent samplings, with error bars representing standard deviations. B RMSE for field and coupling parameters as a function of the number of neurons NN, with the number of trials fixed at $R=550$. Averages and standard deviations are computed over 10 independent samplings. C Average computation time for different numbers of neurons $N$ and trials $R=55,100,300,550$, with standard deviations indicated by error bars. Computation was performed on a Dell PowerEdge R750 server with two Intel Xeon 2.4 GHz CPUs (76 cores / 152 threads).
@@
We evaluated the performance of the proposed state-space kinetic Ising model in terms of estimation accuracy and computational time while varying dataset size and population (See Methods for parameter generation).
@@
Estimation Error To assess estimation accuracy, we computed the root mean squared error (RMSE) between the true parameters ${\theta}_{1:T}$ and the estimated parameters $\theta_{1:t|T}$ for both field and coupling parameters. Specifically, RMSEs were computed separately for the field parameter $\theta_{i,t}$ and the coupling parameter $\theta_{ij,t}$, then averaged over time bins. The means over 10 independent samplings are shown in Figures 3A and B with the standard deviations represented by error bars.
@@
For a fixed number of neurons ($N=80$), RMSEs for both field and coupling parameters decreased as the number of trials $R$ increased (Fig. 3A), demonstrating improved estimation accuracy with more data. Conversely, when the number of trials was fixed at $R=550$, RMSEs exhibited different trends depending on the parameter type. The RMSE for the field parameter increased with $N$, reflecting the challenges of estimating field parameters in larger networks with limited data. However, RMSEs for coupling parameters remained stable across different neuron numbers in this simulation (Fig. 3B).
Computational Time We analyzed computation time for model fitting. The assumption of independent state evolution for individual neurons (Eq. 3) significantly reduces computational complexity by enabling independent calculations for filtering, smoothing, and parameter optimization per neuron, which can be further accelerated through parallel updates. However, heterogeneous couplings increase computational cost.
@@
Figure 3C illustrates the computation time required to complete the EM algorithm for different numbers of neurons $N$ and trials $R$. The results indicate that estimation with $N=80$ and $R=550$ trials can be completed in approximately one hour, making it feasible for practical data analysis. Nevertheless, computation time scales with both $N$ and $R$, highlighting the necessity for further optimization to enable large-scale analysis. One potential improvement is replacing the current filtering method, which employs exact Newton-Raphson optimization for maximum a posteriori (MAP) estimation, with quasi-Newton or mean-field approximations, as demonstrated in equilibrium state-space Ising models [48].
Estimating entropy flow of simulation data
Fig 4: Comparison of entropy flow estimation methods. Entropy flows estimated by four different methods: Sampling method using true parameters ${\boldsymbol{\theta}}_{t}$ (blue); Sampling method using estimated parameters ${\boldsymbol{\theta}}_{t|T}$ (green); Mean-field method using true parameters ${\boldsymbol{\theta}}_{t}$ (black); Mean-field method using estimated parameters $\boldsymbol{\theta}_{t|T}$ (red).
@@
In this section, we assess the proposed mean-field approximation method for estimating entropy flow. As in the previous section, we generated spike samples from time-dependent parameters $\boldsymbol{\theta}_{t}$ sampled from Gaussian processes. All simulations were conducted with $N=80$, $T=75$, and $R=550$ trials. We then estimated the time-dependent field and coupling parameters from the data. Using the posterior mean $\boldsymbol{\theta}_{t|T}$, we obtained the mean-field approximation of the time-dependent entropy flow (Eqs. 13 and 14). The red line in Fig. 4 represents the entropy flow calculated using the mean-field approximation with the learned parameters.
@@
To verify the consistency of the estimated entropy flow, we calculated the entropy flow using a sampling-based method to compute the expectation over the two-step trajectories (blue line). This approach involves repeatedly running the kinetic Ising model (Eq. 1) using the true parameters to sample spike time series data. This process was performed $n_{s}=10,000$ times to empirically estimate the joint distribution $p(\mathbf{x}_{t},\mathbf{x}_{t-1})$. Using this empirical distribution, we obtained a sample estimate of the entropy flow as follows
$\hat{\sigma}_{t}^{\rm flow}=\frac{1}{n_{s}}\sum^{n_{s}}_{s=1}\log\frac{p\left(\mathbf{x}^{s}_{t}\mid\mathbf{x}^{s}_{t-1}\right)}{p\left(\mathbf{x}^{s}_{t-1}\mid\mathbf{x}^{s}_{t}\right)}.$ (21)
@@
This sampling estimation using the true parameters serves as the baseline.
@@
The mean-field estimation of the entropy flow (red) follows the trajectory of the baseline sampling estimation using the true parameters (blue). The result confirms the proposed method is applicable for entropy flow analysis while ensuring computational feasibility. The slight discrepancy between the two lines is due to the errors in estimating the time-dependent parameters and/or the mean-field approximation (in addition to sampling fluctuation inherent to the sampling method). To separate these effects, we estimated the entropy flow by the mean-field approximation using the true parameters $\theta_{t}$ used for the data generation (black). This estimation was deviated from the baseline sampling estimation. In contrast, the sampling method using estimated parameters (green) did not significantly differ from the baseline. Thus, the mean-field approximation introduced errors rather than inaccuracies in parameter estimation. These results indicate that further enhancement of the estimation accuracy is possible by refining the mean-field method.
Model limitations
@@
Fig 5: Analysis of the model miss-specification. A The population spike count histogram of $N=30$ neurons following the shifted-geometric model with a sparseness parameter $f=20$ and $\tau=0.8$ (empirical distribution obtained by the Gibbs sampling in blue circle; theoretical probabilities in green). The yellow line represents a distribution obtained from the state-space kinetic Ising model fitted to the Gibbs sampling data. B The activation function of the shifted-geometric model with $f=20$ and $\tau=0.8$ (green) and that of the kinetic Ising model (yellow) using the average of the fitted field and coupling parameters.
@@
We end the simulation analysis by acknowledging that assumptions of the kinetic Ising framework, in particular pairwise couplings and conditional independence, represent simplifications that may not faithfully capture neural population dynamics. To demonstrate this, we fitted the kinetic Ising model to population activity, using a neuronal population model, called alternating-shrinking higher-order interaction model, that accounts for deviations from the logistic activation function of individual neurons and exhibits higher-order interactions [58].
@@
In this model, homogeneous binary population activity was generated using an exponential-family distribution with interactions of all orders (Eq. 66 in Methods). The model was designed so that the spike-count histogram of the population exhibits sparse yet widespread characteristics (Fig. 5A, green), consistent with empirical data. We performed Gibbs sampling from this distribution (blue circle), which correspond to the dynamics of recurrent networks with an extended activation function.
@@
When the state-space kinetic Ising model was fitted to these activities, it failed to reproduce the observed spike-count histogram (Fig. 5A, yellow). One reason is its restriction to pairwise interactions, which prevents it from capturing higher-order dependencies. Reproducing widespread spike-count histograms in large populations is known to require interactions of all orders [59]. By contrast, the pairwise model concentrates probability mass on only up to two points in the limit of large $N$, often overestimating the tail because it neglects the higher-order interactions that generate sparse, heavy-tailed distributions. The mismatch in model architectures is also apparent in their activation functions (Fig.5B). The alternating-shrinking HOI model exhibits a supra-linear activation function due to the nonlinear integration of synaptic inputs (Eq. 74 in Methods). In contrast, the kinetic Ising model employs the classical logistic activation function with a linear sum of synaptic inputs (Eq. 1).
@@
Application to mouse V1 neurons under active and passive exposures to visual stimuli
@@
In this study, we analyzed the “Allen Brain Observatory: Visual Behavior Neuropixels” dataset provided by the Allen Institute for Brain Science, which contains large-scale recordings of neural spiking activity of mouse brains during the visual change detection task (See [60, 61, 62] for the analyses using this data set). The task is designed to analyze the effect of novelty and familiarity of the stimulus on neural responses. One of two image sets (G and H) were presented to animals at the training/habituation and recording sessions with different orders. The G and H image sets contain 8 natural images. We analyzed the recordings of 37 mice available from the Allen dataset, which were exposed to stimulus G in the recording sessions (either day 1 or 2) whereas the same stimulus G was used in the training and habituation sessions prior to the recording sessions (i.e., the case in which G is familiar).
@@
The neural activities were recorded under two distinct conditions, in which the mice were either actively or passively performing the task under the same set of images. The active condition involved the mice performing a go/no-go change detection task, where they earned a water reward upon detection of a change in the visual stimulus, measured by licking behavior. Each of the 8 stimuli was presented for $250$ ms, followed by a $500$ ms interstimulus interval (gray screen), repeating for one hour while mice actively engaged in the task for reward. In contrast, the passive condition involved replaying the same visual stimuli used in the active condition but without providing any rewards or access to the lick port. In this study, we analyzed recordings with images labeled im036_r, im012_r, and im115_r, which are used in the training session and classified as ‘Familiar,’ and compared neural responses under the active and passive conditions. We used all presentations of the images equally and treated one presentation as a trial.
@@
We selected neurons in the V1 area for analysis. For each mouse, we analyzed simultaneous activities of neurons with length $750$ ms after the onset of the image im036_r, which resulted in $R$ trials. While the number of trials varies across mice, the mean trial number was $R=566$ with $356$ and $652$ as the minimum and maximum number of trials.
Estimating time-dependent parameters of the kinetic Ising model for mouse V1 data: an example
We constructed binary sequences using a 10 ms bin, which results in $T=75$ time bins. Figure 6A (Left) shows the probability of a spiking event within a bin averaged over neurons at each time under the active and passive conditions (population-average spike rate) from an exemplary mouse ($574078$). Here we focus on the analysis of im036_r. The overall temporal profiles were similar across the active and passive conditions. In both conditions, the population exhibited higher mean spike rates during the stimulus presentation period (0-$250$ ms) than the post-stimulus period (250250-$750$ ms). However, their magnitudes differed significantly. The passive condition (blue) showed consistently higher spiking probabilities than the active condition (red) throughout the stimulus and post-stimulus periods. In agreement with the population-average spike-rate dynamics, time-averaged spike rates of individual neurons exhibited a sparser distribution during the active condition compared to the passive condition (Fig. 6A Right). These results were consistently observed across all mice with a few exceptions (S1 Fig and S2 Fig). We compared the average mean activity and sparsity of the distributions under the two conditions for all 37 mice (S3 Fig). Sparsity of a non-negative firing rate distribution was quantified by the coefficient of variation (CV) [63]. The V1 neurons exhibited diminished, sparser firing rate distributions in the active condition than in the passive condition, as confirmed by the reduced mean spike rates ($p=1.556\times 10^{-8}$, Wilcoxon signed-rank test) and increased CVs ($p=8.35\times 10^{-8}$, Wilcoxon signed-rank test).
@@
Fig 6: Estimated neural dynamics under the active and passive conditions for mouse $574078$. A Spike-rate dynamics and distributions. (Left) Spiking probabilities averaged across neurons and trials. Time points ($t=5,25,35,50$) used in panel C. (Right) Spike-rate distributions of all recorded neurons. B Estimated time-dependent parameters $\boldsymbol{\theta}_{t|T}$ of the kinetic Ising model under active (top) and passive (bottom) conditions. Each row shows estimates of the field $\theta_{i,t}$ and three coupling $\theta_{ij,t}$ panels for $i=1,2,80$. Solid lines and shaded areas indicate MAP estimates and pointwise 6868% intervals ($\pm 1$ SD). For each $i$, we display the top-55 significant pairs within the analysis window (bins 21217575) whose intervals exclude zero at least once; among them, the smallest $j$ labels are shown. C Estimated couplings at $t=5,25,35,50$ under the active (top) and passive (bottom) conditions. Neurons are indexed 118080 in the original ordering.
@@
We then applied the state-space kinetic Ising model to the binary activities of these neurons. For this goal, we selected the top $N=80$ neurons with the highest spike rates. The estimated dynamics of the field and coupling parameters exhibited variations in both active and passive conditions (Fig. 6B). Notably, the field parameters $\theta_{i,t}$ (the first column) follow the dynamics of the mean spike rate of the population with significant fluctuations. On the contrary, the dynamics of the coupling parameters $\theta_{ij,t}$ exhibited smoother transitions. To clarify the dynamics of the couplings, we show them in the matrix form at specific time points, $t=5,25,35,50$ (Fig. 6C; timing marked in black lines in Fig. 6A left). The neurons are indexed in the ascending order of the average firing rates. The top and bottom rows show the results of the active and passive conditions, respectively. Coupling strength is indicated by color, with red and blue representing positive and negative values, respectively. The results show that (i) the couplings exhibit significant variations with positive and negative values; (ii) the variations are stronger in the active condition than in the passive condition; (iii) the diagonal components of the couplings (self-correlations) mostly display negative correlations.
@@
To corroborate the above observations, we performed the same analysis on the trial-shuffled data (S4 Fig). Trial-shuffled data reveal bias and variance in estimation under the assumption of neuronal independence. The result shows a significant reduction in the magnitude and variability of the couplings whereas self-couplings remained unchanged (note that the self-coupling remains after trial-shuffling). However, non-zero couplings persisted with stronger variations in the active condition than in the passive condition, reflecting sampling fluctuations due to the lower firing rates in the active condition. These findings indicate that the parameters observed in Fig.6B,C include estimation noise, necessitating statistical analyses to confirm their significance.

### v3

@@ Results
The term $r\left(g_{i}+z\sqrt{\Delta_{i}}\right)-m_{i}$ quantifies how the neuron’s activity rate deviates from its long-term average, while $z\sqrt{\Delta_{i}}$ represents the fluctuations of the input it receives. The steady-state mean-field solution thus provides an intuitive view of entropy flow as a measure of a neuron’s causal responsiveness to input fluctuations – a quantity that captures the correlation underlying Hebbian plasticity in neural systems. However, this equation also clarifies that the approximation depends mainly on the magnitudes of the field and coupling parameters and is thus insensitive to the detailed coupling structure. It should therefore be applied with caution when the degree of coupling asymmetry is the primary determinant of the strength of entropy flow.
Simulation: Estimating the model parameters
@@
We begin by testing the proposed method by estimating the time-dependent parameters of a kinetic Ising model consisting of two simulated neurons (Fig. 1A). Figure 1B shows the spike data generated using Eq. 1 with the number of bins, $T=400$, and the number of trials, $L=200$. The time-dependent parameters ${\boldsymbol{\theta}}_{1:T}$ used to generate binary data were sampled from Gaussian processes (See Methods).
@@
The EM algorithm was applied to this spike data until the log marginal likelihood converged (Fig. 1C). Figure 1D shows the components of the optimized hyperparameter matrices, $\mathbf{Q}^{i}(i=1,2)$. Figures 1E show the MAP estimates of the time-dependent fields $\theta_{i,t}$ and couplings $\theta_{ij,t}$ under the optimized hyperparameters (solid lines) with 95% credible intervals (shaded areas). The results confirm that the method uncovers the underlying time-dependent parameters (black dashed lines) used to generate the data.
@@
Figure 2: The application of the state-space kinetic Ising model to 12 simulated neurons. A Simulated spike data for the first, 100th, and last trial out of 200 trials. The vertical axis shows the number of neurons, and the horizontal axis represents the number of bins. B Estimated coupling parameters $\boldsymbol{\theta}_{t|T}$ (solid lines), for all neurons and time bins ($i=1,2,\ldots,12$, $t=1,\ldots,T$). Shaded areas indicate 95% credible intervals, and dashed lines denote the true parameter values used to generate the data. These plots show only the couplings that are significantly deviated from zero: The couplings whose 95% credible interval contains 0 in all bins were excluded. For clarity, only five such significant incoming couplings from other neurons are shown in each panel. C Scatter plots comparing the true coupling parameters ${\boldsymbol{\theta}}_{t}$ with the estimated values $\boldsymbol{\theta}_{t|T}$ at time $t=10,20,\ldots,60$. The black line is a diagonal line.
@@
Next, we applied the state-space kinetic Ising model to a network of 12 simulated neurons to estimate the time-varying field and coupling parameters between neurons. Figure 2A presents the spike data generated using the observation model with the number of bins set to $T=75$ and the number of trials $L=200$. Data generation and model estimation procedures follow the two-neuron case above. Figure 2B shows the estimated time-varying coupling parameters $\theta_{ij,t}$ for each neuron. In Fig. 2C, we compare the estimated coupling parameters $\boldsymbol{\theta}_{t|T}$ with the true values $\boldsymbol{\theta}_{t}$ at representative time points ($t=10,20,\ldots,60$). The scatter plot shows agreement between the true and estimated values, with most points aligning closely along the diagonal line, indicating that the model captured the underlying dynamics of the coupling parameters. These results confirm that the proposed state-space kinetic Ising model can reliably estimate time-varying coupling parameters in a network of simulated neurons.
Simulation: Estimation error and computational time
@@
Figure 3: Estimation error and computational time. A Root mean squared error (RMSE) of the field and coupling parameter estimation as a function of trials $L$, with the number of neurons fixed at $N=80$. Results are averaged over 10 independent samples, with error bars representing standard deviations. B RMSE of the field and coupling parameters as a function of the number of neurons NN, with the number of trials fixed at $L=550$. Averages and standard deviations are computed over 10 independent samples. C Average computation time for different numbers of neurons $N$ and trials $L=55,100,300,550$, with error bars indicating standard deviations. Computation was performed on a Dell PowerEdge R750 server with two Intel Xeon 2.4 GHz CPUs (76 cores / 152 threads).
@@
We evaluated the performance of the proposed state-space kinetic Ising model in terms of estimation accuracy and computational time, varying dataset and population sizes (See Methods for parameter generation).
@@
Estimation error: To assess estimation error, we computed the root mean squared error (RMSE) between the true parameters ${\boldsymbol{\theta}}_{t}$ and the estimated parameters $\boldsymbol{\theta}_{t|T}$ for both field and coupling parameters. Namely, RMSEs were computed separately for the field parameter $\theta_{i,t}$ and the coupling parameter $\theta_{ij,t}$, then averaged over time bins. The means over 10 independent samplings are shown in Figures 3A and B with the standard deviations represented by error bars.
@@
For a fixed number of neurons ($N=80$), RMSEs for both field and coupling parameters decreased as the number of trials $L$ increased (Fig. 3A), demonstrating improved estimation accuracy with more data. Conversely, when the number of trials was fixed at $L=550$, RMSEs exhibited different trends depending on the parameter type. The RMSE for the field parameter increased with $N$, imposing the challenges of estimating field parameters in larger networks with limited data. The RMSE for coupling parameters remained stable across different neuron numbers in this simulation (Fig. 3B).
@@
Computational time: We analyzed the computation time for model fitting. Figure 3C illustrates the computation time required to complete the EM algorithm for different numbers of neurons $N$ and trials $L$. The results indicate that estimation with $N=80$ and $L=550$ trials can be completed in approximately one hour, making it feasible for practical data analysis. Nevertheless, computation time scales with both $N$ and $L$, highlighting the necessity for further optimization to enable large-scale analysis. The assumption of independent state evolution for individual neurons (Eq. 3) significantly reduces computational complexity by enabling independent calculations for filtering, smoothing, and parameter optimization per neuron, which can be further accelerated through parallel updates. Another potential improvement is replacing the current filtering method, which employs exact Newton-Raphson optimization for maximum a posteriori (MAP) estimation, with quasi-Newton or mean-field approximations, as demonstrated in equilibrium state-space Ising models [48].
Simulation: Estimating entropy flow
@@
In this section, we assess the proposed mean-field approximation method for estimating entropy flow. As in the previous section, we generated spike samples from time-dependent parameters $\boldsymbol{\theta}_{1:T}$ sampled from Gaussian processes. All simulations were conducted with $N=80$, $T=75$, and $L=550$ trials. We then estimated the time-dependent field and coupling parameters from the data. Using the posterior mean $\boldsymbol{\theta}_{t|T}$, we obtained the mean-field approximation of the time-dependent entropy flow (Eq. 12, using Eqs. 13 and 14). The solid red line in Fig. 4 represents the entropy flow calculated using the mean-field approximation with the learned parameters.
Figure 4: Comparison of entropy flow estimation methods. Entropy flows estimated using four different approaches: Sampling method with true parameters ${\boldsymbol{\theta}}_{t}$ (solid black); sampling method with estimated parameters ${\boldsymbol{\theta}}_{t|T}$ (dashed green); mean-field method with true parameters ${\boldsymbol{\theta}}_{t}$ (dashed blue); and mean-field method with estimated parameters $\boldsymbol{\theta}_{t|T}$ (solid red).
@@
To verify the consistency of the estimated entropy flow, we calculated the entropy flow using a sampling-based method to compute the expectation over the two-step trajectories (solid black). This approach involves repeatedly running the kinetic Ising model (Eq. 1) using the true parameters to sample binary spike sequences. This process was performed $n_{s}=10,000$ times to empirically estimate the joint distribution $p(\mathbf{x}_{t},\mathbf{x}_{t-1})$. Using this empirical distribution, we obtained a sample estimate of the entropy flow as follows:
$\hat{\sigma}_{t}^{\rm flow}=\frac{1}{n_{s}}\sum^{n_{s}}_{s=1}\log\frac{p\left(\mathbf{x}^{s}_{t}{|}\mathbf{x}^{s}_{t-1}\right)}{p\left(\mathbf{x}^{s}_{t-1}{|}\mathbf{x}^{s}_{t}\right)},$ (21)
@@
where $\mathbf{x}^{s}_{t}$ denotes the $s$-th sample at time $t$. This sampling estimation using the true parameters serves as the baseline.
@@
The mean-field estimation of the entropy flow (solid red) follows the trajectory of the baseline sampling estimation using the true parameters (solid black). The result confirms that the proposed method is applicable for entropy flow analysis while ensuring computational feasibility. The slight discrepancy between the two lines is due to the errors in estimating the time-dependent parameters and/or the mean-field approximation (in addition to sampling fluctuation inherent to the sampling method). To separate these effects, we estimated the entropy flow by the mean-field approximation using the true parameters $\boldsymbol{\theta}_{1:T}$ used for the data generation (dashed blue). This estimation deviated from the baseline sampling estimation. In contrast, the sampling method using estimated parameters (dashed green) did not significantly differ from the baseline. Thus, the discrepancy arose from the mean-field approximation, rather than from inaccuracies in parameter estimation. These results suggest that refining the mean-field method could further improve the accuracy of entropy flow estimation.
Simulation: Model limitations
@@
Figure 5: Analysis of the model misspecification. A The population spike count histogram of $N=30$ neurons following the shifted-geometric model with a sparseness parameter $f=20$ and $\tau=0.8$ (empirical distribution obtained by the Gibbs sampling in blue circle; theoretical probabilities in green). The yellow line represents a distribution obtained from the state-space kinetic Ising model fitted to the Gibbs sampling data. B The activation function of the shifted-geometric model with $f=20$ and $\tau=0.8$ (green) and that of the kinetic Ising model (yellow) using the average of the fitted field and coupling parameters.
@@
We end the simulation analysis by acknowledging that assumptions of the kinetic Ising framework, in particular pairwise couplings and conditional independence, represent simplifications that may not faithfully capture neural population dynamics. To demonstrate this, we fitted the kinetic Ising model to population activity, using a neuronal population model called the alternating-shrinking higher-order interaction model, which accounts for deviations from the logistic activation function of individual neurons and exhibits higher-order interactions [58].
@@
In this model, homogeneous binary population activity was generated using an exponential-family distribution with interactions of all orders (Eq. 63 in Methods). The model was designed so that the spike-count histogram of the population exhibits sparse yet widespread characteristics (Fig. 5A, green), consistent with empirical data. We performed Gibbs sampling from this distribution (blue circle), which corresponds to the dynamics of recurrent networks with an extended activation function (See Methods).
@@
When the state-space kinetic Ising model was fitted to these activities, it failed to reproduce the observed spike-count histogram (Fig. 5A, yellow). One reason is its restriction to pairwise interactions, which prevents it from capturing higher-order dependencies. Reproducing widespread spike-count histograms in large populations is known to require interactions of all orders [59]. By contrast, the pairwise model concentrates probability mass on only up to two points in the limit of large $N$, often overestimating the tail because it neglects the higher-order interactions that generate sparse, heavy-tailed distributions. The mismatch in model architectures is also apparent in their activation functions (Fig.5B). The alternating-shrinking higher-order interaction model exhibits a supra-linear activation function due to the nonlinear integration of synaptic inputs (Eq. 71 in Methods). In contrast, the kinetic Ising model employs the classical logistic activation function with a linear sum of synaptic inputs (Eq. 1).
@@
Mouse V1 neurons: Experimental design and data description
@@
In this study, we analyzed the Allen Brain Observatory: Visual Behavior Neuropixels dataset provided by the Allen Institute for Brain Science, which contains large-scale recordings of neural spiking activity of mouse brains during the visual change detection task (See [60, 61, 62] for analyses using this data set). The task is designed to analyze the effect of novelty and familiarity of the stimulus on neural responses. One of two image sets (G and H) was presented to animals at the training/habituation and recording sessions with different orders. The G and H image sets contain 8 natural images. We analyzed the recordings of 37 mice available from the Allen dataset, which were exposed to stimulus G in the recording sessions (either day 1 or 2) whereas the same stimulus G was used in the training and habituation sessions prior to the recording sessions (i.e., the case in which G is familiar).
@@
The neural activities were recorded under two distinct conditions, in which the mice were either actively or passively performing the task under the same set of images. The active condition involved the mice performing a go/no-go change detection task, where they earned a water reward upon detection of a change in the visual stimulus, measured by licking behavior. Each of the 8 stimuli was presented for $250$ ms, followed by a $500$ ms interstimulus interval (gray screen), repeating for one hour while mice actively engaged in the task for reward. In contrast, the passive condition involved replaying the same visual stimuli used in the active condition but without providing any rewards or access to the lick port. In this study, we analyzed recordings with images labeled im036_r, im012_r, and im115_r, which are used in the training session and classified as Familiar, and compared neural responses under the active and passive conditions. We used all presentations of the images equally and treated one presentation as a trial.
@@
We selected neurons in the V1 area for analysis. For each mouse, we analyzed the simultaneous activity of neurons during a $750$-ms period following the image onset. Although the number of trials varies across mice, the mean trial count was $566$ with $356$ and $652$ as the minimum and maximum number of trials, respectively, for the case of an image im036_r.
Mouse V1 neurons: An exemplary result from a single mouse
@@
Figure 6: Estimated neural dynamics under the active and passive conditions for mouse $574078$. A Spike-rate dynamics and distributions. (Left) Spike-rate averaged across neurons and trials. (Right) Spike-rate distributions of all recorded neurons. B Smoothed time-dependent parameters $\boldsymbol{\theta}_{t|T}$ of the kinetic Ising model for the active (top) and passive (bottom) conditions. The first column shows the field $\theta_{i,t}$ (one trace per neuron), and the next three columns show the incoming couplings $\theta_{ij,t}$ for $i=1,2,80$. Solid lines are MAP estimates and shaded areas indicate $\pm 1$ SD (i.e., 68% credible bands) computed from the diagonal of the posterior covariance. For each $i$, couplings were first screened within the analysis window (bins 2121-$75$) and retained if their credible interval excluded zero at least once in the window (self-couplings excluded). From the retained set, we display the first five couplings per $i$, ordered by ascending $j$ label for readability. C Estimated couplings at $t=5,25,35,50$ under the active (top) and passive (bottom) conditions.
We constructed binary sequences using a 10 ms bin, which resulted in $T=75$ time bins. Here, we focused on the analysis of im036_r. Figure 6A (Left) shows the spike-rate averaged over neurons at each time under the active and passive conditions (population-average spike rate) from an exemplary mouse ($574078$). The overall temporal profiles were similar across the active and passive conditions. In both conditions, the population exhibited higher mean spike rates during the stimulus presentation period (0-$250$ ms) than the post-stimulus period (250250-$750$ ms). However, their magnitudes significantly differed across the conditions. The passive condition (blue) showed consistently higher spiking probabilities than the active condition (red) throughout the stimulus and post-stimulus periods. In agreement with the population-average spike-rate dynamics, time-averaged spike rates of individual neurons exhibited a sparser distribution during the active condition compared to the passive condition (Fig. 6A Right).
@@
We then applied the state-space kinetic Ising model to the binary activities of these neurons. For this goal, we selected the top $N=80$ neurons with the highest spike rates. The estimated dynamics of the field and coupling parameters exhibited variations in both active and passive conditions (Fig. 6B). Notably, the field parameters $\theta_{i,t}$ (the first column) follow the dynamics of the mean spike rate of the population with significant fluctuations. On the contrary, the dynamics of the coupling parameters $\theta_{ij,t}$ exhibited smoother transitions. To clarify the dynamics of the couplings, we show them in the matrix form at specific time points, $t=5,25,35,50$ (Fig. 6C). The neurons are indexed in the ascending order of the average firing rates. The top and bottom rows show the results of the active and passive conditions, respectively. Coupling strength is indicated by graded color, with red and blue representing positive and negative values, respectively. The results show that (i) the couplings exhibit significant variations with positive and negative values; (ii) the variations are stronger in the active condition than in the passive condition; (iii) the diagonal components of the couplings (self-correlations) mostly display negative correlations.
@@
To corroborate the above observations, we performed the same analysis on the trial-shuffled data (Supplementary Fig. S1). The analysis of trial-shuffled data reveals bias and variance in estimation under the assumption of neuronal independence. The result shows a significant reduction in the magnitude and variability of the couplings, whereas self-couplings remained unchanged (note that the self-coupling remains after trial-shuffling). However, non-zero couplings persisted with stronger variations in the active condition than in the passive condition, reflecting sampling fluctuations due to the lower firing rates in the active condition. These findings indicate that the parameters observed in Fig.6B,C include estimation noise, necessitating statistical analyses to confirm their significance.
Mouse V1 neurons: Population analysis across mice